Louis walks and jogs to school each day. He averages 5 km/h walking and 9 km/h jogging. The distance from home to school is 8 km and he makes the trip in 1 hour. How far does he jog in a trip?
Answer by richwmiller(17219)

w+j=1,
5w+9j=8
jogging j = 3/4, walking w = 1/4
9*3/4=27/4
6.25 km jogging
5*1/4=1.25 km walking